# Transfer Learning Between U.S. Presidential Elections: How Should We Learn From A 2020 Ad Campaign To Inform 2024 Ad Campaigns?

### v1

@@ Abstract
For the 2024 U.S. presidential election, would negative, digital ads against Donald Trump impact voter turnout in Pennsylvania (PA), a key “tipping point” state? The gold standard to address this question, a randomized experiment where voters get randomized to different ads, yields unbiased estimates of the ad effect, but is very expensive. Instead, we propose a less-than-ideal, but significantly cheaper and likely faster framework based on transfer learning, where we transfer knowledge from a past ad experiment in 2020 to evaluate ads for 2024. A key component of our framework is a sensitivity analysis that quantifies the unobservable differences between past and future elections, which can be calibrated in a data-driven manner. We propose two estimators of the 2024 ad effect: a simple regression estimator with bootstrap, which we recommend for practitioners in this field, and an estimator based on the efficient influence function for broader applications. Using our framework, we estimate the effect of running a negative, digital ad campaign against Trump on voter turnout in PA for the 2024 election. Our findings indicate effect heterogeneity across counties of PA and among important subgroups stratified by gender, urbanicity, and education attainment.
Keywords: Causal inference; Exponential tilting; Generalizability; Sensitivity analysis; Transportability
@@ 1.1 Motivation: Learning From Past Ad Campaigns
During every election, campaigns run ads to influence the political narrative and voter turnout and in recent years, evaluated ad effectiveness through experiments. For example, in the most recent, 2020 U.S. presidential election, Aggarwal et al., (2023) conducted a large-scale, randomized experiment among $1,999,282$ registered voters from Pennsylvania (PA), Wisconsin (WI), Michigan (MI), North Carolina (NC) and Arizona (AZ). They found that a negative, digital ad campaign against Donald Trump during the 2020 U.S. presidential election was ineffective in changing voter turnout. For other examples, see Gerber et al., (2011) and Kalla and Broockman, (2018).
@@
The main question we address in this paper is inspired by Aggarwal et al., (2023): similar to 2020, would negative digital ads against Trump be ineffective in changing voter turnout for the 2024 U.S. presidential election? In both 2020 and 2024 elections, Trump is the nominee for the Republican party and digital anti-Trump ads have been used throughout. But, as Aggarwal et al., (2023) cautioned, their null results may not generalize to elections that are less exceptional than 2020, notably when COVID-19 emerged. In 2024, voters face new issues including women’s rights, inflation, immigration, the Russo-Ukraine War and the Israel-Hamas War (Pew Research Center,, 2024; Ipsos Core Political,, 2024). Also, after the attack on the U.S. Capitol Building on January 6, 2021, a recent study (Loving and Smith,, 2024) found that a significant number of Republican voters left their party and did not return. In short, there are unmeasurable and measurable changes between 2020 and 2024 in terms of electoral contexts and voter demographics.
@@
The ideal approach to answer the paper’s question is to run a randomized experiment in 2024, similar to the one in Aggarwal et al., (2023). However, ad campaigns are expensive; Aggarwal et al., (2023)’s experiment costed \$8.9 million U.S. dollars (USD). More recently, one super political action committee for the Democratic Party, which was characterized as “an ad-making laboratory…testing thousands of messages, social media posts and ads in the 2024 race, ranking them in order of effectiveness and approving only those that resonate with voters”, has already spent \$450 million USD for the 2024 election (Schleifer and Goldmacher,, 2024).
@@
Our approach to answer the question is less than ideal, but arguably significantly cheaper and likely quicker. Specifically, we use transfer learning with sensitivity analysis to “transfer” knowledge from the existing 2020 experiment by Aggarwal et al., (2023) while accounting for changes in context and voter demographics between 2020 and 2024. More formally, transfer learning uses the overlapping measurements about voters in 2020 and 2024 (e.g., gender, age group, party affiliation, voting history) to “adjust” for measurable differences between the two elections. We then use parameters from a sensitivity analysis to quantify any unmeasured differences between 2020 and 2024 (e.g., changes in electoral context) and propose an intuitive, data-driven procedure to calibrate/benchmark the magnitude of these parameters based on sample splitting and an idea from design sensitivity (Rosenbaum,, 2004, 2020). For our data analysis, this calibration procedure assesses whether the unmeasured differences between 2020 and 2024 are larger compared to those between “blue-collar” states and non-blue-collar states in 2020; see Section 6.1 for details.
@@
With our framework, we estimate the effect of negative, digital ads against Trump on voter turnout for the 2024 U.S. presidential election. We focus on roughly $4.9$ million registered voters in Pennsylvania (PA) as our target population. PA is not only the largest swing state in terms of electoral votes, but also the “tipping point state” for the 2024 U.S. presidential election (e.g., FitzGerald, (2024); FiveThirtyEight, (2024)). We present a county-by-county analysis of the ad effect and a subgroup analysis among $20$ politically important subgroups. We find that if 2020 and 2024 elections are similar with respect to some of the voters’ demographics (see Section 6.1), the negative digital ad campaign against Trump would decrease voter turnout in Fulton county for the 2024 election. Interestingly, Fulton county had the largest share of votes for Trump in 2020 among all 67 counties in PA (i.e., 85.41% votes for Trump). But, the ads are estimated to remain ineffective in all other counties of PA for the 2024 election. Also, if there are slight, unmeasured differences between 2020 and 2024, the ad campaign against Trump can change voter turnout in 55 out of 67 counties in PA; see Section 6 for details on interpreting sensitivity parameters. Comparing our results with the results of the 2020 election, our analysis suggests that the negative ads in 2024 impact mostly Trump-leaning counties in terms of depressing voter turnout. In the subgroup analysis, we find that anti-Trump ads can decrease voter turnout among female voters living in rural areas with low college education and increase voter turnout among non-female voters living in urban areas with high college education.
@@ 1.2 Related Work and Contributions
Our work builds upon several works on generalizing or transporting treatment effects from a source population (i.e., voters in the 2020 election) to a target population (i.e., voters in the 2024 election) under a sensitivity analysis framework (Nguyen et al.,, 2017; Nie et al.,, 2021; Colnet et al.,, 2021; Dahabreh et al.,, 2023; Zeng et al.,, 2023; Duong et al.,, 2023; Ek and Zachariah,, 2023; Huang, 2024b, ). Specifically, we work under the exponential tilting sensitivity model (Robins et al.,, 2000), which has been used in works on generalizability and transportability (Franks et al.,, 2020; Dahabreh et al.,, 2020; Scharfstein et al.,, 2021; Dahabreh et al.,, 2022), and make the following contributions.
@@
    We allow the source data (i.e., the 2020 data from Aggarwal et al., (2023)’s experiment)) to have more covariates than the target data (i.e., the 2024 PA voter registration data). Not only was this the case in our own data analysis, but this setting is also common when the source data is derived from a randomized experiment where detailed information about the study units is collected. Zeng et al., (2023) also considered this setup for a similar reason, but focused on efficient and minimax estimation. See also Supplement C.3 of Jin and Rothenhäusler, (2024) who considered conditional inference under this setup.
@@
    We propose a simple regression estimator with nonparametric, percentile bootstrap to estimate the treatment effect in the target population. This is our recommended procedure for practitioners evaluating ads, or, more generally, evaluating treatments/policies for a target population where there are both measurable and unmeasurable differences between the source and target populations. We recommend this procedure because of its simplicity, theoretically attractive properties (e.g., consistency, asymptotic normality), and because the estimator based on the efficient influence function (EIF) lacks the “rate double robustness” property; see below. Also, while qualitatively suggested in several works in this area, we formally show one theoretically correct approach to conduct bootstrap-based inference for transfer learning; see Section 4.1.
@@
    We also propose an estimator based on the EIF. This result extends the novel results in Zeng et al., (2023) to the setting where sensitivity parameters are present. While this estimator is more widely applicable than that in (b), it is more complex, necessitates estimating four nuisance functions and is not doubly robust; see Section 4.2
@@
    For either procedure (b) or (c), we propose a simple calibration procedure to generate a plausible, “reference” magnitude of the sensitivity parameter. Often in sensitivity analysis, practitioners are unsure whether a sensitive parameter is small or large as the parameter represents unobserved differences that cannot be adjusted by measured confounders. Unlike existing methods for calibration (e.g., Hsu and Small, (2013); Cinelli and Hazlett, (2020); Huang, 2024b ) based on omitting a measured covariate, our calibration procedure uses the same covariates for calibration and sensitivity analysis and is inspired by a clever idea underlying design sensitivity (Rosenbaum,, 2004, 2020) and sample splitting. Specifically, we create a data-driven “favorable” situation (Rosenbaum,, 2020, Chapter 15) by splitting the source data in Aggarwal et al., (2023) and restricting the sensitivity parameters to values that are plausible in the partitioned data; see Section 5.
While the listed contributions are directly motivated from the statistical challenges in our data analysis, we believe they can be meaningful in other contexts, notably in generalizing the results of a randomized trial to a target population. More broadly, we hope our analysis pipeline centered on sensitivity analysis with transfer learning is useful to practitioners who want a simple, but theoretically valid approach for generalization or transportation tasks.
@@ 2.1 Setup: Observed Data
Suppose we collect $n_{s}$ i.i.d. samples from a source population and for each study unit $i\in{\cal I}_{s}=\{1,\ldots,n_{s}\}$, we observe the following data:
@@
The variable ${\bf X}_{i}\in{\cal X}$ is the pre-treatment covariate (e.g., voter demographics), $A_{i}$ is the binary treatment indicator (e.g., assigned to ad campaign against Trump or not), $Y_{i}$ is the binary outcome (e.g., voted or not), and $S_{i}$ indicates whether the study unit is from the source sample (i.e., $S_{i}=1$) or not (i.e., $S_{i}=0$). In our data analysis, the source data is from Aggarwal et al., (2023) where there were $n_{s}=1,999,282$ registered voters from the five states: PA, WI, MI, NC, and AZ. We also collect $n_{t}$ i.i.d. samples from the target population and for each study unit $i\in{\cal I}_{t}=\{n_{s}+1,\ldots,n_{s}+n_{t}=n\}$, we observe the following data [^1] For notational convenience, we overload the notation ${\bf O}_{i}$ to represent the observed data from unit $i$. If the data is from the source, ${\bf O}_{i}=({\bf X}_{i}$, $A_{i}$, $Y_{i},S_{i}=1)$ and if the data is from the target, ${\bf O}_{i}=({\bf V}_{i},S_{i}=0)$.:
@@
The variable ${\bf V}_{i}\in{\cal V}\subseteq{\cal X}$ is a subset of the covariates in $\cal{X}$. In our data analysis, the target data is registered voters from PA’s voter registration database and ${\bf V}_{i}$ is voter $i$’s demographic information in the database (e.g., age group, gender, party affiliation, voting history). Because ${\bf V}_{i}$ is present in both the source and the target data, we refer to it as a shared covariate. Figure 2.1 summarizes our data setup.
@@
We make some remarks about the setup. First, if the covariates are discrete, some modeling assumptions about the outcome regression or the propensity score in Sections 4.1 and 4.2 are automatically satisfied. In our analysis, all covariates were discrete, including the voter’s age, which was measured by Aggarwal et al., (2023) as an age group. Second, we allow ${\cal V}\subseteq{\cal X}$ because, as far as we are aware of, there is no publicly available dataset of the 2024 voter population that measured the same attributes about voters as the source data from 2020. In general, we find that if the source population is from a randomized controlled trial, the covariates from it (i.e., ${\bf X}_{i}$) are richer than those from the target population (i.e., ${\bf V}_{i}$); see Zeng et al., (2023) who echoed a similar sentiment. Third, while we focus on binary outcomes $Y_{i}$ due to our data analysis, our framework generalizes to a continuous outcome; see Section A of the Appendix. Fourth, similar to other works in transfer learning, we assume that the units in the source and the target data are independent and sampled from an infinite population in order to derive asymptotic properties of our estimators below. But, this may be a poor approximation in some settings; see Jin and Rothenhäusler, (2024) for detailed discussions. Section 7 discusses this issue in the context of our data analysis.
2.2 Setup: Causal Estimands and Nuisance Parameters
@@
$\theta=\theta_{1}-\theta_{0},\text{ where }\theta_{a}=\mathbb{E}(Y^{(a)}_{i}\mid S_{i}=0)\text{ and }a\in\{0,1\}.$
In our data analysis, $\theta$ is the average effect of a digital ad campaign against Trump on voter turnout among registered PA voters in 2024. We remark that for a binary outcome, other measures of treatment effects are possible, such as the risk ratio $\theta_{1}/\theta_{0}$ and the odds ratio $[\theta_{1}/(1-\theta_{1})][\theta_{0}/(1-\theta_{0})]^{-1}$. While we focus on mean differences (i.e. $\theta_{1}-\theta_{0}$) like Aggarwal et al., (2023), our results are derived for $\theta_{a}$ and thus, can be extended to cover the risk ratio and the odds ratio; see Ye et al., (2023) for an example.
We define the following functions, often referred to as nuisance parameters.
@@
  • The ratio of probability densities of ${\bf V}_{i}$ between the target and the source populations: $w({\bf v})=p_{{\bf V}\mid S=0}({\bf v}\mid S_{i}=0)/p_{{\bf V}\mid S=1}({\bf v}\mid S_{i}=1)$ where $p_{{\bf V}\mid S=s}(\cdot)$ denotes the conditional density of ${\bf V}_{i}$ given $S_{i}=s$ and ${\bf v}\in{\cal V}$.
We conclude by defining the following notations for order and convergences. For two real sequences of numbers $b_{n}$ and $d_{n}$, we denote $b_{n}=O(b_{n})$ if $|b_{n}|\leq C|d_{n}|$ for a constant $C$ and denote $b_{n}\asymp d_{n}$ if $b_{n}=O(d_{n})$ and $d_{n}=O(b_{n})$. We use $\to_{p}$ to denote convergence in probability and $\to_{d}$ to denote convergence in distribution. For a random variable $Z_{n}$ and a real sequence of numbers $b_{n}$ we denote $Z_{n}=o_{p}(b_{n})$ if $Z_{n}/b_{n}\to_{p}0$. For a measurable and integrable function $f(\cdot)$, we denote its $L_{2}$ norm with respect to $\mathbb{P}$ by $\lVert f({\bf O}_{i})\rVert=\sqrt{\mathbb{E}\{f^{2}({\bf O}_{i})\}}$.

### v2

@@ Abstract
For the 2024 U.S. presidential election, would negative, digital ads against Donald Trump impact voter turnout in Pennsylvania (PA), a key “tipping point” state? The gold standard to address this question, a randomized experiment where voters get randomized to different ads, yields unbiased estimates of the ad effect, but is very expensive. Instead, we propose a less-than-ideal, but significantly cheaper and faster framework based on transfer learning, where we transfer knowledge from a past ad experiment in 2020 to evaluate ads for 2024. A key component of our framework is a sensitivity analysis that quantifies the unobservable differences between 2020 and 2024 elections, where sensitivity parameters can be calibrated in a data-driven manner. We propose two estimators of the 2024 ad effect: a simple regression estimator with bootstrap, which we recommend for practitioners in this field, and an estimator based on the efficient influence function for broader applications. Using our framework, we estimate the effect of running a negative, digital ad campaign against Trump on voter turnout in PA for the 2024 election. Our findings indicate effect heterogeneity across counties of PA and among important subgroups stratified by gender, urbanicity, and education attainment.
Keywords: Causal inference; Generalizability; Sensitivity analysis; Transportability; Exponential tilting
@@ 1.1 Motivation: Learning from past ad campaigns
In recent years, political campaigns have used randomized experiments to evaluated political ads (e.g., Gerber et al., (2011); Kalla and Broockman, (2018); Aggarwal et al., (2023)). For example, in the 2020 U.S. presidential election, Aggarwal et al., (2023) conducted a large-scale, randomized controlled trial (RCT) among $1,999,282$ registered voters from Pennsylvania (PA), Wisconsin (WI), Michigan (MI), North Carolina (NC) and Arizona (AZ). They found that a negative, digital ad campaign against President Donald Trump during the 2020 U.S. presidential election was ineffective in changing voter turnout.
@@
The main empirical question we address in this paper is as follows: similar to 2020, would negative digital ads against Trump remain ineffective in changing voter turnout for the 2024 U.S. presidential election? In both 2020 and 2024 elections, Trump was the nominee for the Republican party and digital anti-Trump ads were used extensively. But, as Aggarwal et al., (2023) pointed out, 2020 was an exceptional election due to COVID-19 and their null results may not generalize to less exceptional elections. In particular, compared to 2020, voters faced new issues in 2024 including women’s rights, inflation, the Russo-Ukraine War and the Israel-Hamas War (Pew Research Center,, 2024; Ipsos Core Political,, 2024). Also, Loving and Smith, (2024) showed that after the attack on the U.S. Capitol Building on January 6, 2021, some Republican voters left their party, signaling a potential shift in voter demographics between the 2020 and the 2024 elections. In short, there were measurable and unmeasurable differences between 2020 and 2024 in terms of electoral contexts and voter demographics.
@@
The ideal approach to answer the main empirical question is to re-run the randomized experiment by Aggarwal et al., (2023). While this approach can yield unbiased estimates of the ad effect irrespective of the differences between 2020 and 2024, ad campaigns are very expensive. For example, Aggarwal et al., (2023)’s experiment in 2020 costed \$8.9 million U.S. dollars (USD). More recently, one super political action committee for the Democratic Party, which was characterized as “an ad-making laboratory…testing thousands of messages, social media posts and ads in the 2024 race, ranking them in order of effectiveness”, spent \$450 million USD for the 2024 election (Schleifer and Goldmacher,, 2024).
@@
Our approach to answer the empirical question is less than ideal, but significantly cheaper and faster. Specifically, we use transfer learning with sensitivity analysis to “transfer” knowledge from the existing, 2020 experiment by Aggarwal et al., (2023) while accounting for measurable and unmeasurable changes in electoral context and voter demographics between 2020 and 2024. More formally, transfer learning uses the overlapping measurements about voters in 2020 and 2024 (e.g., gender, age group, party affiliation, voting history) to “adjust” for measurable differences between the two elections. We then use parameters from a sensitivity analysis to quantify any unmeasured differences between the elections (e.g., changes in electoral context). We also propose a new, data-driven procedure to calibrate/benchmark the magnitude of the parameters from the sensitivity analysis based on sample splitting and design sensitivity (Rosenbaum,, 2004, 2020).
@@
With our framework, we estimate the effect of negative, digital ads against Trump on voter turnout for the 2024 U.S. presidential election. We focus on roughly $4.9$ million registered voters in Pennsylvania (PA) as our target population. PA is not only the largest swing state in terms of electoral votes, but also the “tipping point state” for the 2024 U.S. presidential election (e.g., FiveThirtyEight, (2024)). We present a county-by-county analysis of the ad effect and a subgroup analysis among $20$ politically important subgroups. In the county-by-county analysis, we find that if 2020 and 2024 elections are similar with respect to some of the voters’ demographics, the negative digital ad campaign against Trump would decrease voter turnout in Fulton county, a heavily Republican-leaning county. But, the ads would remain ineffective in all other counties of PA for the 2024 election. Moreover, if there are slight, unmeasured differences between 2020 and 2024, the ad effects change from being insignificant to significant in 60 counties. In the subgroup analysis, we find that the negative ads can decrease voter turnout among female voters in rural areas with low college education and increase turnout among non-female voters in urban areas with high college education.
@@ 1.2 Related work and contributions
Our work builds upon several works on generalizing or transporting treatment effects from a source population to a target population under a sensitivity analysis framework (Nguyen et al.,, 2017; Colnet et al.,, 2021; Dahabreh et al.,, 2023; Zeng et al.,, 2023; Duong et al.,, 2023; Ek and Zachariah,, 2023; Huang, 2024b, ). Specifically, we work under the exponential tilting sensitivity model (Robins et al.,, 2000), which has been used in works on generalizability and transportability (Dahabreh et al.,, 2022), and make the following new contributions.
@@
    We allow the source data to have more covariates than the target data. Not only was this the case in our own data analysis, but this setting is common when the source data is derived from a randomized experiment where detailed information about the study units is collected. Zeng et al., (2023) also considered this setup for a similar reason, but focused on efficient and minimax estimation.
@@
    We propose a simple regression estimator with nonparametric, percentile bootstrap to estimate the treatment effect in the target population; see Section 4.1. Notably, while qualitatively suggested by several works in this area, we formally show one theoretically correct approach to conduct bootstrap-based inference for transfer learning. We recommend this analysis pipeline for practitioners because of its simplicity, theoretically attractive properties (e.g., consistency, asymptotic normality), and the estimator based on the efficient influence function (EIF) is not doubly robust; see below.
@@
    We also propose an estimator based on the EIF. This result extends the novel results in Zeng et al., (2023) to the setting where sensitivity parameters are present. While this estimator is more widely applicable than that in (b), it is more complex, requires estimating four nuisance functions and is not doubly robust; see Section 4.2
@@
    For either procedure (b) or (c), we propose a simple calibration procedure to generate interpretable, “reference” magnitudes of the sensitivity parameter. Unlike existing methods for calibration based on omitting a measured covariate (e.g., Hsu and Small, (2013); Cinelli and Hazlett, (2020); Ek and Zachariah, (2023); Huang, 2024b ), our calibration procedure uses the same covariates for both calibration and sensitivity analysis. The calibration procedure is inspired by a clever idea underlying design sensitivity (Rosenbaum,, 2004) and sample splitting where we create a data-driven “favorable” situation (Rosenbaum,, 2020, Chapter 15) by splitting the source data in a particular way; see Section 5.
While the listed contributions are directly motivated from the statistical challenges in our data analysis, we believe they can be meaningful in other contexts, notably in generalizing the results of a randomized trial to a target population with mis-matching covariates and unmeasurable differences between the populations. More broadly, we hope our analysis pipeline centered on sensitivity analysis with transfer learning is useful to practitioners who want a simple, theoretically valid approach for generalization or transportation tasks.
@@ 2.1 Setup: Observed data
Suppose we collect $n_{s}$ independent and identically distributed (i.i.d.) samples from a source population. For each study unit $i\in{\cal I}_{s}=\{1,\ldots,n_{s}\}$ in the source data, we observe the following:
@@
The variable ${\bf X}_{i}\in{\cal X}$ is the pre-treatment covariate (e.g., voter demographics), $A_{i}$ is the binary treatment indicator (e.g., assigned to ad campaign against Trump or not), $Y_{i}$ is the binary outcome (e.g., voted or not), and $S_{i}$ indicates whether unit $i$ is from the source sample (i.e., $S_{i}=1$) or not (i.e., $S_{i}=0$). In our data analysis, the source data is from Aggarwal et al., (2023). Independently, we also collect $n_{t}$ i.i.d. samples from the target population and for each study unit $i\in{\cal I}_{t}=\{n_{s}+1,\ldots,n_{s}+n_{t}=n\}$ from the target data, we observe the following [^1] For notational convenience, we overload the notation ${\bf O}_{i}$ to represent the observed data from unit $i$. If the data is from the source, ${\bf O}_{i}=({\bf X}_{i}$, $A_{i}$, $Y_{i},S_{i}=1)$ and if the data is from the target, ${\bf O}_{i}=({\bf V}_{i},S_{i}=0)$.:
@@
The variable ${\bf V}_{i}\in{\cal V}\subseteq{\cal X}$ is a subset of the covariates in $\cal{X}$. In our data analysis, the target data consists of registered voters from PA’s voter registration database and ${\bf V}_{i}$ is voter $i$’s demographic information in the database (e.g., age group, gender, party affiliation, voting history). Because ${\bf V}_{i}$ is present in both the source and the target data, we refer to it as the shared covariate. Figure 2.1 summarizes our data setup.
@@
We make some remarks about the setup. First, if the covariates are discrete, some modeling assumptions about the outcome regression or the propensity score in Sections 4.1 and 4.2 are automatically satisfied. In our data analysis, all covariates were discrete. Second, we allow ${\cal V}\subseteq{\cal X}$ because, as far as we are aware of, there is no publicly available dataset of the 2024 voter population that measured the same attributes about voters as the source data from 2020. In general, we find that if the source population is from a randomized controlled trial, the covariates from it (i.e., ${\bf X}_{i}$) are richer than those from the target population (i.e., ${\bf V}_{i}$); see Zeng et al., (2023) who echoed a similar sentiment. Third, while we focus on binary outcomes $Y_{i}$ due to our data analysis, our framework generalizes to a continuous outcome; see Section A of the Appendix. Fourth, similar to other works on transfer learning, we assume that the units in the source and the target data are independent and sampled from an infinite population in order to derive asymptotic properties of our estimators below. But, this may lead to conservative inference in some settings (Jin and Rothenhäusler,, 2024) and Section 7 discusses this issue in the context of our data analysis.
2.2 Setup: Causal estimands and nuisance functions
@@
$\theta=\theta_{1}-\theta_{0},\text{ where }\theta_{a}=\mathbb{E}\left[Y^{(a)}_{i}\mid S_{i}=0\right]\text{ and }a\in\{0,1\}.$
In our data analysis, $\theta$ is the average effect of a digital ad campaign against Trump on voter turnout in 2024 among registered PA voters. We remark that for a binary outcome, other measures of treatment effects are possible, such as the risk ratio $\theta_{1}/\theta_{0}$ and the odds ratio $[\theta_{1}/(1-\theta_{1})][\theta_{0}/(1-\theta_{0})]^{-1}$. While we focus on mean differences (i.e. $\theta_{1}-\theta_{0}$) like Aggarwal et al., (2023), our results are derived for $\theta_{a}$ and thus, can be extended to cover the risk ratio and the odds ratio; see Ye et al., (2023) for an example.
We define the following functions, often referred to as nuisance functions.
@@
  • The ratio of probability densities of ${\bf V}_{i}$ between the two populations: $w({\bf v})=p_{{\bf V}_{i}\mid S_{i}=0}({\bf v}\mid S_{i}=0)/p_{{\bf V}_{i}\mid S_{i}=1}({\bf v}\mid S_{i}=1)$ where $p_{{\bf V}_{i}\mid S_{i}=s}(\cdot)$ is the conditional density of ${\bf V}_{i}$ given $S_{i}=s$, $s=0,1$.
We conclude by defining the following notations for order and convergences. For two real sequences of numbers $b_{n}$ and $d_{n}$, we denote $b_{n}=O(b_{n})$ if $|b_{n}|\leq C|d_{n}|$ for a constant $C$ and denote $b_{n}\asymp d_{n}$ if $b_{n}=O(d_{n})$ and $d_{n}=O(b_{n})$. We use $\to_{p}$ to denote convergence in probability and $\to_{d}$ to denote convergence in distribution. For a sequence of random variables $Z_{n}$ and a real sequence of numbers $b_{n}$, we denote $Z_{n}=o_{p}(b_{n})$ if $Z_{n}/b_{n}\to_{p}0$. For a measurable and integrable function $f(\cdot)$, we denote its $L_{2}$ norm by $\lVert f({\bf O}_{i})\rVert=\sqrt{\mathbb{E}\{f^{2}({\bf O}_{i})\}}$.